# On the connection between Bochner’s theorem on positive definite maps and Choi theorem on complete positivity

### v1

@@ II.2 CJKS Isomorphism and completely positive maps
where $\tau$ represent transposition with respect to the usual canonical basis of $\mathcal{H}_{1}$ and $\mathbb{I}$ represents the identity operator on $\mathcal{H}_{2}$. The Eq. (14) will be referred as the Choi inversion formula throughout the paper.
A map $\Phi:B(\mathcal{H}_{1})\rightarrow B(\mathcal{H}_{2})$ is called positive if it maps positive elements of $B(\mathcal{H}_{1})$ to the positive elements of $B(\mathcal{H}_{2})$. Now using $\Phi$, one can construct a map $\mathrm{id}_{k}\otimes\Phi:M_{k}(\mathbb{C})\otimes B(\mathcal{H}_{1})\rightarrow M_{k}(\mathbb{C})\otimes B(\mathcal{H}_{2})$, where $M_{k}(\mathbb{C})$ is the space of $k\times k$ matrices whose entries are complex numbers.
@@
The map $\Phi:B(\mathcal{H}_{1})\rightarrow{B}(\mathcal{H}_{2})$ is called $k$-positive if $\mathrm{id}_{k}\otimes\Phi:M_{k}(\mathbb{C})\otimes B(H_{1})\rightarrow M_{k}(\mathbb{C})\otimes B(H_{2})$ is positive. If $\Phi$ is $k$-positive for all $k\in\mathbb{N}$, then $\Phi$ is called completely positive.
From the definition it is obvious that a complete positive map is a positive map, however the converse is not always true. The transposition map on matrices is an example of a map which is positive but not completely positive. The following theorem provides necessary and sufficient condition for a map to be completely positive.
Theorem L (Choi Theorem).
[5, 6, 1, 7, 8]. A map $\Phi:B(\mathcal{H}_{1})\rightarrow B(\mathcal{H}_{2})$ is completely positive if and only if the Choi matrix $C_{\Phi}=\sum_{{i,j}=1}^{n}e_{ij}\otimes\Phi(e_{ij})$ is positive semidefinite.
II.3 Space of Super-maps
We consider the Hilbert spaces $\mathcal{H}_{1}$, $\mathcal{H}_{2}$, $\mathcal{H}_{3}$ and $\mathcal{H}_{4}$ to be finite dimensional. Let $B(\mathcal{H}_{i})$ represents the space of bounded linear operators on the Hilbert space $\mathcal{H}_{i}$, for $i=1,2,3,4$. We represent by $L(B(\mathcal{H}_{i}),B(\mathcal{H}_{j}))$ the space of linear operators from $B(\mathcal{H}_{i})$ to $B(\mathcal{H}_{j})$ and by $L[L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2})),L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))]$ we represent the space of linear operators from $L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2}))$ to $L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))$. A linear map $\Theta:L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2}))\rightarrow L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))$ is called a super-map.
II.3.1 Choi-type theorem for Super-maps
@@
The spaces $L\left[L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2})),L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))\right]$ and $L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2}))\otimes L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))$ are isomorphic as vector spaces under the following identification
$\Theta\leftrightarrow\sum_{ijkl}\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}\otimes\Theta(\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}).$ (16)
It should be noted that the object $\Lambda_{\Theta}:=\sum_{ijkl}\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}\otimes\Theta(\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl})\in L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2}))\otimes L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))$ has the mathematical structure similar to Choi matrices (see Ref.[10])
III Main Results
@@
Using the fact that in the group multiplication table, along a row or column each element of the group appears once and only once, the above sum can be rewritten as
@@
where $(f_{1}*f_{2})(g_{k}):=\sum_{i=1}^{|G|}f_{1}(g_{i})f_{2}(g_{i}^{-1}g_{k})$ is the convolution of $f_{1}$ and $f_{2}$. This shows that the convolution algebra $L(\mathbb{C}_{0}[G],\mathbb{C})$ is isomorphic to the group algebra $\mathbb{C}_{0}[G]$. The above fact can be generalized as follows:
@@
See Appendix (B) for the proof. Now we generalize the above theorem in the context of contracted semigroup algebra. We begin by defining convolution in the space $L(\mathbb{C}_{0}[S],\mathcal{A})$ of linear maps from $\mathbb{C}_{0}[S]$ to an arbitrary algebra $\mathcal{A}$.
@@
$\mathrm{\Phi}*\mathrm{\Phi^{\prime}}(s_{k}):=\begin{cases}\sum_{i,j}\mathrm{\Phi}(s_{i})\mathrm{\Phi^{\prime}}(s_{j})&\text{if}\ s_{k}\in\mathrm{Im}(\cdot),\\ 0&\text{if}\ s_{k}\notin\mathrm{Im}(\cdot).\end{cases}$ (19)
for all $s\in S\setminus z$. Where, the sum is over the pair $(i,j)$ such that $s_{i}\cdot s_{j}=s_{k}$ and $\mathrm{Im}(\cdot)$ represents the image of the multiplication map $\cdot:S\times S\rightarrow S$ of the semigroup $S$.
@@
Let $S$ be a semigroup with zero element $z$. We denote the contracted algebra of $S$ by $\mathbb{C}_{0}[S]$. Then the convolution algebra $L(\mathbb{C}_{0}[S],\mathcal{A})$ is isomorphic to $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ with the convolution product being defined by Eq. (19).
@@
where the summation inside the bracket of the first term is over those pair of elements $(s_{p},s_{q})$ such that $s_{p}s_{q}=s_{k}\neq 0$ and the summation inside the bracket of the second term is over the pair $(u,v)$ such that $s_{u}\cdot s_{v}=0$. As the second term vanishes, we have
@@
Let us now consider a $d$-dimensional Hilbert space $\mathcal{H}$. The set $S:=\{e_{ij}\}_{i,j=1}^{d}\cup\{0\}\subset\mathcal{B}(\mathcal{H})$ with “0” being the zero matrix, is a semigroup under the matrix multiplication:
@@
Then the contracted algebra $\mathbb{C}_{0}[S]$ is isomorphic to the full algebra $\mathcal{B}(\mathcal{H})$ [3, 11]. Hence, we can view the matrix algebra $\mathcal{B}(\mathcal{H})$ as the contracted semigroup algebra of the semigroup $\{e_{ij}\}_{i,j=1}^{d}\cup\{0\}$. In the specific case when we consider the space $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ of linear maps from $\mathcal{B}(\mathcal{H}_{1})$ to $\mathcal{B}(\mathcal{H}_{2})$ then Theorem (1) becomes the following:
The convolution algebra $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ and the algebra $\mathcal{B}(\mathcal{H}_{1})\otimes\mathcal{B}(\mathcal{H}_{2})$ are isomorphic, with the isomorphism being the Choi-Jamiołkowski isomorphism $\Phi\mapsto\sum_{i,j}e_{ij}\otimes\Phi(e_{ij})$ and the convolution product being
$\Phi*\Phi^{\prime}(e_{ij}):=\sum_{k}\Phi(e_{ik})\Phi^{\prime}(e_{kj})$ (26)
@@
Recalling that $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ and $L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))$ are convolution algebras with the convolution being defined as in Eq. (26), and taking in to account Proposition (1), we define convolution on $L\left[L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2})),L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))\right]$ as follows:
@@
In this setting Theorem (1) becomes the following:
@@
A super-map $\mathrm{\Theta}:L\left(\mathcal{B}\left(\mathcal{H}_{1}\right),\mathcal{B}\left(\mathcal{H}_{2}\right)\right)\rightarrow L\left(\mathcal{B}\left(\mathcal{H}_{3}\right),\mathcal{B}\left(\mathcal{H}_{4}\right)\right)$ induces a map $T:\mathcal{B}\left(\mathcal{H}_{1}\right)\otimes\mathcal{B}\left(\mathcal{H}_{2}\right)\rightarrow\mathcal{B}\left(\mathcal{H}_{3}\right)\otimes\mathcal{B}\left(\mathcal{H}_{4}\right)$ known as “representing map” [12] at the level of Choi matrices and is given by the following,
@@
We now prove that the map $\mathrm{\Theta}\leftrightarrow T$ preserves the respective convolutions. Let $T\leftrightarrow\Theta_{1}\star\Theta_{2}$. Then we want to show that $T=T_{1}*T_{2}$:
@@
With the above observation, the convolution algebras $L[L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2})),L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))]$ and $L[\mathcal{B}\left(\mathcal{H}_{1}\right)\otimes\mathcal{B}\left(\mathcal{H}_{2}\right),\mathcal{B}\left(\mathcal{H}_{3}\right)\otimes\mathcal{B}\left(\mathcal{H}_{4}\right)]$ are isomorphic.
III.1 Fourier transform of maps
Let $G$ be a finite group and $f:G\rightarrow\mathbb{C}$ be a complex-valued function on $G$. Let $\rho:G\rightarrow M_{d_{\rho}}(\mathbb{C})$ be a representation of the group $G$ of degree $d_{\rho}$. The Fourier transform of the function $f$ w.r.t the representation $\rho$ is given by:

### v2

@@ 2.2 Choi-Jamiołkowski Isomorphism and Completely Positive Maps
where $\tau$ denotes transposition with respect to the standard basis of $\mathcal{H}_{1}$, and $\mathbb{I}$ denotes the identity operator on $\mathcal{H}_{2}$. Equation 21 will be referred to as the Choi inversion formula throughout the paper.
A linear map $\Phi:\mathcal{B}(\mathcal{H}_{1})\rightarrow\mathcal{B}(\mathcal{H}_{2})$ is called positive if it maps positive elements of $\mathcal{B}(\mathcal{H}_{1})$ to positive elements of $\mathcal{B}(\mathcal{H}_{2})$. Given a linear map $\Phi$, one can construct a map
$\mathrm{id}_{k}\otimes\Phi:M_{k}(\mathbb{C})\otimes\mathcal{B}(\mathcal{H}_{1})\rightarrow M_{k}(\mathbb{C})\otimes\mathcal{B}(\mathcal{H}_{2}),$
where $M_{k}(\mathbb{C})$ denotes the algebra of $k\times k$ complex matrices.
@@
A linear map $\Phi:\mathcal{B}(\mathcal{H}_{1})\rightarrow\mathcal{B}(\mathcal{H}_{2})$ is called $k$-positive if $\mathrm{id}_{k}\otimes\Phi:M_{k}(\mathbb{C})\otimes\mathcal{B}(H_{1})\rightarrow M_{k}(\mathbb{C})\otimes\mathcal{B}(H_{2})$ is positive. If $\Phi$ is $k$-positive for all $k\in\mathbb{N}$, then $\Phi$ is called completely positive.
By definition, every complete positive map is positive. However, the converse is not always true. A standard example is the transposition map on matrices, which is positive but fails to be completely positive. The following theorem provides a necessary and sufficient condition for a map to be completely positive.
Theorem L (Choi’s Theorem [29, 30, 15, 31, 32]).
A linear map $\Phi:\mathcal{B}(\mathcal{H}_{1})\rightarrow\mathcal{B}(\mathcal{H}_{2})$ is completely positive if and only if its Choi matrix $C_{\Phi}=\sum_{{i,j}=1}^{n}e_{ij}\otimes\Phi(e_{ij})$ is positive semidefinite.
2.2.1 Complete Order Isomorphism
A linear map $\Phi:M_{n}(\mathbb{C})\to M_{m}(\mathbb{C})$ is called a complete order isomorphism if both $\Phi$ and $\Phi^{-1}$ are completely positive. An operator $X\in M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})$ is said to satisfy CP vs. positivity correspondence if the complete positivity of a linear map $\Phi:M_{n}(\mathbb{C})\to M_{m}(\mathbb{C})$ is equivalent to the positivity of $(\mathrm{id}\otimes\Phi)X$ for all $m\in\mathbb{N}$ [33].
Theorem M (Kye, Paulsen [33, 34]).
Let $\sigma:M_{n}(\mathbb{C})\to M_{n}(\mathbb{C})$ be a linear map. Then the Choi matrix $C_{\sigma}$ satisfies the CP vs. positivity correspondence if and only if $\sigma$ is a complete order isomorphism.
2.3 Space of Super-maps
Let $\mathcal{H}_{i}$ be finite dimensional Hilbert spaces for $i\!\in\!\{1,2,3,4\}$, and let $\mathcal{B}(\mathcal{H}_{i})$ denote the algebra of bounded linear operators on $\mathcal{H}_{i}$. We denote by
$L(\mathcal{B}(\mathcal{H}_{i}),\mathcal{B}(\mathcal{H}_{j}))$
the space of linear operators from $\mathcal{B}(\mathcal{H}_{i})$ to $\mathcal{B}(\mathcal{H}_{j})$, and by
$L[L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2})),L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))]$
we denote the space of linear operators from $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ to $L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))$. A linear map $\Theta:L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))\rightarrow L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))$ is called a super-map.
@@ 2.3 Space of Super-maps
Let us denote by $\mathbb{CP}[1\rightarrow 2]$ and $\mathbb{CP}[3\rightarrow 4]$ the convex cone of completely positive maps of $L(\mathcal{B}(H_{1}),\mathcal{B}(H_{2}))$ and $L(\mathcal{B}(H_{3}),\mathcal{B}(H_{4}))$ respectively.
Definition.
A super-map $\Theta:L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))\rightarrow L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))$ is called CP-preserving if it maps CP maps in $\mathbb{CP}[1\rightarrow 2]$ to the CP maps in $\mathbb{CP}[3\rightarrow 4]$, i.e., $\Theta(\mathbb{CP}[1\rightarrow 2])\subseteq\mathbb{CP}[3\rightarrow 4]$.
Definition.
A super-map $\Theta:L(\,\mathcal{B}(\mathcal{H}_{1}),\,\mathcal{B}(\mathcal{H}_{2}))\rightarrow L(\,\mathcal{B}(\mathcal{H}_{3}),\,\mathcal{B}(\mathcal{H}_{4}))$ is called completely CP preserving if the map $\mathrm{id}\otimes\Theta:L(\,\mathcal{B}(K_{1}),\,\mathcal{B}(K_{2}))\otimes L(\,\mathcal{B}(\mathcal{H}_{1}),\,\mathcal{B}(\mathcal{H}_{2}))\rightarrow L(\,\mathcal{B}(K_{1}),\,\mathcal{B}(K_{2}))\otimes L(\,\mathcal{B}(\mathcal{H}_{3}),\,\mathcal{B}(\mathcal{H}_{4}))$ is CP-preserving for for all choices of Hilbert spaces $K_{1}$ and $K_{2}$.
The spaces $L\left[L(\,\mathcal{B}(\mathcal{H}_{1}),\,\mathcal{B}(\mathcal{H}_{2})),L(\,\mathcal{B}(\mathcal{H}_{3}),\,\mathcal{B}(\mathcal{H}_{4}))\right]$ and $L(\,\mathcal{B}(\mathcal{H}_{1}),\,\mathcal{B}(\mathcal{H}_{2}))\otimes L(\,\mathcal{B}(\mathcal{H}_{3}),\,\mathcal{B}(\mathcal{H}_{4}))$ are isomorphic as vector spaces under the identification
$\Theta\leftrightarrow\Lambda_{\Theta}:=\sum_{ijkl}\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}\otimes\Theta(\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}).$ (23)
It should be noted that the object $\Lambda_{\Theta}$ has the mathematical structure similar to Choi matrices.
Theorem ([23, 24, 18]).
The super-map
$\Theta:L(B(H_{1}),B(H_{2}))\rightarrow L(B(H_{3}),B(H_{4}))$
is completely CP-preserving if and only if
$\Lambda_{\Theta}:=\sum_{ijkl}\mathcal{E}^{H_{1}\rightarrow H_{2}}_{ijkl}\otimes\Theta(\mathcal{E}^{H_{1}\rightarrow H_{2}}_{ijkl})$
is CP.
3 Main Results
In this section, we present the formal statements and proofs of the main results outlined in the introduction. We begin with the structural properties of the convolution algebra on $\mathbb{C}_{0}[S]$ and proceed to the Fourier inversion and Plancherel formulas for finite inverse semigroups, followed by our Bochner-type theorem and its reduction to the Choi’s theorem on complete positivity.
@@ 3 Main Results
Using the fact that in group multiplication table, each element of $G$ appears exactly once in every row and every column, the above sum can be rewritten as
@@
where
$(f_{1}*f_{2})(g_{k}):=\sum_{i=1}^{|G|}f_{1}(g_{i})f_{2}(g_{i}^{-1}g_{k})$
is the convolution of $f_{1}$ and $f_{2}$. This shows that the convolution algebra $L(\mathbb{C}_{0}[G],\mathbb{C})$ is isomorphic to the group algebra $\mathbb{C}_{0}[G]$. This observation admits a natural generalization as follows:
@@
See Appendix B for the proof. Now we generalize the above theorem in the context of contracted semigroup algebra. We begin by defining convolution in the space $L(\mathbb{C}_{0}[S],\mathcal{A})$ of linear maps from $\mathbb{C}_{0}[S]$ to an arbitrary algebra $\mathcal{A}$.
@@
$\mathrm{\Phi}*\mathrm{\Phi^{\prime}}(s_{k}):=\begin{cases}\sum_{i,j}\mathrm{\Phi}(s_{i})\mathrm{\Phi^{\prime}}(s_{j})&\text{if}\ s_{k}\in\mathrm{Im}(\cdot),\\ 0&\text{if}\ s_{k}\notin\mathrm{Im}(\cdot),\end{cases}$ (26)
for all $s_{k}\in S\setminus\{z\}$, where the sum is over the pairs $(i,j)$ such that $s_{i}\cdot s_{j}=s_{k}$, and $\mathrm{Im}(\cdot)$ denotes the image of the multiplication map $\cdot:S\times S\rightarrow S$ of the semigroup $S$.
@@
Let $S$ be a semigroup with zero element $z$ and denote the contracted algebra of $S$ by $\mathbb{C}_{0}[S]$. Then the convolution algebra $L(\mathbb{C}_{0}[S],\mathcal{A})$ is isomorphic to $\mathbb{C}_{0}[S]\otimes\mathcal{A}$, where the convolution product is defined by Eq. (26).
@@
where the summation inside the bracket of the first term is over those pair of elements $(s_{p},s_{q})$ such that $s_{p}s_{q}=s_{k}\neq 0$ and the summation inside the bracket of the second term is over the pairs $(u,v)$ such that $s_{u}\cdot s_{v}=0$. As the second term vanishes, we have
@@
Let us now consider a $d$-dimensional Hilbert space $\mathcal{H}$, and denote the matrix units of $\mathcal{B(H)}$ by $\{e_{ij}\}_{i,j=1}^{d}$. The set $S:=\{e_{ij}\}_{i,j=1}^{d}\cup\{0\}\subset\mathcal{B}(\mathcal{H})$ with “0” being the zero matrix, is a semigroup under the matrix multiplication:
@@
Then the contracted algebra $\mathbb{C}_{0}[S]$ is isomorphic to the full algebra $\mathcal{B}(\mathcal{H})$ [26, 25]. Hence, we can view the matrix algebra $\mathcal{B}(\mathcal{H})$ as the contracted semigroup algebra of the semigroup $\{e_{ij}\}_{i,j=1}^{d}\cup\{0\}$. In the specific case where we consider the space $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ of linear maps from $\mathcal{B}(\mathcal{H}_{1})$ to $\mathcal{B}(\mathcal{H}_{2})$, Theorem 1 becomes the following:
The convolution algebra $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ and the algebra $\mathcal{B}(\mathcal{H}_{1})\otimes\mathcal{B}(\mathcal{H}_{2})$ are isomorphic, with the isomorphism being the Choi-Jamiołkowski isomorphism
$\Phi\mapsto\sum_{i,j}e_{ij}\otimes\Phi(e_{ij}),$
and the convolution product being
$\Phi*\Phi^{\prime}(e_{ij}):=\sum_{k}\Phi(e_{ik})\Phi^{\prime}(e_{kj}),$ (33)
@@
Recalling that $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ and $L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))$ are convolution algebras with the convolution being defined as in Eq. (33), and taking in to account Proposition 1, we define convolution on $L\left[L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2})),L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))\right]$ as follows:
@@
In this setting Theorem 1 becomes the following:
@@
A super-map
$\mathrm{\Theta}:L\left(\mathcal{B}\left(\mathcal{H}_{1}\right),\mathcal{B}\left(\mathcal{H}_{2}\right)\right)\rightarrow L\left(\mathcal{B}\left(\mathcal{H}_{3}\right),\mathcal{B}\left(\mathcal{H}_{4}\right)\right)$
induces a map
$T:\mathcal{B}\left(\mathcal{H}_{1}\right)\otimes\mathcal{B}\left(\mathcal{H}_{2}\right)\rightarrow\mathcal{B}\left(\mathcal{H}_{3}\right)\otimes\mathcal{B}\left(\mathcal{H}_{4}\right)$
at the level of Choi matrices, known as the “representing map” [23] of $\Theta$ and it is given by
@@
We now prove that the map $\mathrm{\Theta}\leftrightarrow T$ preserves the respective convolutions. Let $T\leftrightarrow\Theta_{1}\star\Theta_{2}$. Then we aim to show that $T=T_{1}*T_{2}$:
@@
With the above observation, it follows that the convolution algebras $L[L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2})),L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))]$ and $L[\mathcal{B}\left(\mathcal{H}_{1}\right)\otimes\mathcal{B}\left(\mathcal{H}_{2}\right),\mathcal{B}\left(\mathcal{H}_{3}\right)\otimes\mathcal{B}\left(\mathcal{H}_{4}\right)]$ are isomorphic.
3.1 Fourier Transform of Maps
Let $G$ be a finite group, and $f:G\rightarrow\mathbb{C}$ be a complex-valued function on $G$. Let $\rho:G\rightarrow M_{d_{\rho}}(\mathbb{C})$ be a representation of the group $G$ of degree $d_{\rho}$. The Fourier transform of the function $f$ with respect to the representation $\rho$ is given by: